# Interleaved dual-species arrays of single atoms using a passive optical element and one trapping laser

### v1

@@ 1 Introduction
Arrays of trapped neutral atoms are being rapidly developed for quantum sensing, communication, and computing applications[1]. Scaling up to large numbers of trapped atoms in well-ordered arrays is necessary for large-scale computation and other quantum-information applications. Large arrays have been achieved using sophisticated optical systems comprising high-power lasers and active beam-shaping components such as acousto-optic deflectors (AODs) [2], spatial light modulators (SLMs) [3, 4], and digital mirror devices (DMDs) [5, 6]. New developments utilize interleaved arrays of two different atomic species for error correction and noise reduction[7, 8, 9, 10] where, for example, one type of atom can be a computation qubit and the other a measurement qubit. Trapping multiple species of atoms simultaneously adds to system complexity, especially when using multiple trapping lasers with different wavelengths and active electro-optical components [8]. System complexity can be reduced by using specially fabricated passive components to generate arrays of optical traps[11, 12, 13].
@@ 5 Trapping Rb and Cs atoms
Atoms were first loaded into a dual-species MOT. The magnetic quadrupole gradient was abruptly turned off and the detuning of the cooling light from resonance was increased to further cool the atoms via polarization gradient cooling (PGC) [21]. During the PGC phase, atoms were stochastically loaded into the optical tweezer (bright) and bottle (dark) traps, where they underwent collisional blockade [22] until at most one atom per trap site survived. The resulting atom arrays were then imaged by pulsing quasi-resonant light from the side of the vacuum cell (shown in Fig. 6(a)) to induce fluorescence at wavelengths 780 nm for Rb and 852 nm for Cs. This fluorescence was collected via a second high-NA objective on the other side of the vacuum cell. A series of dichroic beamsplitters was used to separate the trapping light from the fluorescence and to split the fluorescence light into two separate imaging paths, one for each species, before recombining the images of the two interleaved arrays onto a single electron-multiplying charge-coupled device (EMCCD). The imaging separation of Rb and Cs atoms helps increase the signal to background of each species. The averaged fluorescence image of 5000 loading shots collected by the EMCCD is shown in Fig. 6 (a) with the Rb array of atoms in the lower-left and the Cs array in the upper-right. Each atom-trapping site is spatially separated by about 6 .

### v2

@@ 1 Introduction
Arrays of trapped neutral atoms are being rapidly developed for quantum sensing, communication, and computing applications[1]. Scaling up to large numbers of trapped atoms in well-ordered arrays is necessary for large-scale computation and other quantum-information applications. Large arrays have been achieved using sophisticated optical systems comprising high-power lasers and active beam-shaping components such as acousto-optic deflectors (AODs) [2], spatial light modulators (SLMs) [3, 4], and digital mirror devices (DMDs) [5, 6]. New developments utilize interleaved arrays of two different atomic species for error correction and noise reduction[7, 8, 9, 10] where, for example, one type of atom can be a computation qubit and the other a measurement qubit. Trapping multiple species of atoms simultaneously adds to system complexity, especially when using multiple trapping lasers with different wavelengths and active electro-optical components [8]. System complexity can be reduced by using specially fabricated passive components to generate arrays of optical traps[11, 12, 13, 14, 15].
@@ 5 Trapping Rb and Cs atoms
Atoms were first loaded into a dual-species MOT. The magnetic quadrupole gradient was abruptly turned off and the detuning of the cooling light from resonance was increased to further cool the atoms via polarization gradient cooling (PGC) [23]. During the PGC phase, atoms were stochastically loaded into the optical tweezer (bright) and bottle (dark) traps, where they underwent collisional blockade [24] until at most one atom per trap site survived. The resulting atom arrays were then imaged by pulsing quasi-resonant light from the side of the vacuum cell (shown in Fig. 6(a)) to induce fluorescence at wavelengths 780 nm for Rb and 852 nm for Cs. This fluorescence was collected via a second high-NA objective on the other side of the vacuum cell. A series of dichroic beamsplitters was used to separate the trapping light from the fluorescence and to split the fluorescence light into two separate imaging paths, one for each species, before recombining the images of the two interleaved arrays onto a single electron-multiplying charge-coupled device (EMCCD, Nuvu HNü512). The imaging separation of Rb and Cs atoms helps increase the signal to background of each species. The averaged fluorescence image of 5000 loading shots collected by the EMCCD is shown in Fig. 6 (a) with the Rb array of atoms in the lower-left and the Cs array in the upper-right. Each atom-trapping site is spatially separated by about 6 .